OpenStudy (anonymous):

Two masses of 1Kg and 2Kg are connected by a light inextensilbe atring which passes over a light , frictionless ,fixed pulley.The masses are initially at rest the velocity of the mass at the instant one of the masses has moved up a distance of 7.5 m is (m)

OpenStudy (anonymous):

@ajprincess @sauravshakya @Ishaan94

OpenStudy (anonymous):

20-T = 2a T-10 = a

OpenStudy (ajprincess):

ya. find a.

OpenStudy (anonymous):

3a = 10 a = 10/3

OpenStudy (anonymous):

V^2 - u^2 = 2as V^2 = 2*10/3 * 7.5 = 50

OpenStudy (anonymous):

the answer should be 7

OpenStudy (ajprincess):

v=7.07106 when we round it off v=7.

OpenStudy (anonymous):

oh....ok..

OpenStudy (anonymous):

thxxx

OpenStudy (ajprincess):

yw.:)